A plane wave has equation y=25sin (120t-4x). Find the
(I) amplitude
(II) wavelength
(III) wave velocity
(iv) frequency and period of the wave

(where y and x are in metres, t is in seconds)
1
Expert's answer
2019-02-20T13:49:19-0500

The general equation describing a plane wave looks like:


"y(x, t) = Asin(\omega t - kx),"


here, "A" is the amplitude of the plane wave, "k" is the wavenumber, "\omega" is the angular frequency of the plane wave.

I) As we can see from the equation above, the amplitude of the wave is:


"A = 25 m."


II) We can find the wavelength from the formula:


"k = \frac{2 \pi}{\lambda},"


here, "\lambda" is the wavelength of the wave, "k = 4 \frac{rad}{m}" is the wavenumber.

Then, we get:


"\lambda = \frac{2 \pi}{k} = \frac{2 \pi}{4 \dfrac{rad}{m}} = 1.57 m."


III) We can find the velocity of the wave from the wave speed formula:


"v = f \lambda,"


here, "f" is the frequency of the wave.

We can find the frequency of the wave from the formula:


"\omega = 2 \pi f,""f = \dfrac{\omega}{2 \pi}."

Then, substituting the freuency into the ave speed formula, we can find the velocity of the wave:


"v = f \lambda = \dfrac{\omega \lambda}{2 \pi} = \dfrac{120 \dfrac{rad}{s} \cdot 1.57 m}{2 \pi} = 30 \dfrac{m}{s}."

IV) We can find the frequency of the wave from the formula:


"f = \dfrac{\omega}{2 \pi} = \dfrac{120 \dfrac{rad}{s}}{2 \pi} = 19.1 Hz."


We can find the period of the ave from the formula:


"T = \dfrac{1}{f} = \dfrac{1}{19.1 Hz} = 0.052 s."

Answer:

I) "A = 25 m."

II) "\lambda = 1.57 m."

III) "v = 30 \dfrac{m}{s}."

IV) "f = 19.1 Hz", "T = 0.052 s."